OpenStudy (anonymous):

X^X+X^X=839 what is x?

OpenStudy (anonymous):

do u have options by chance!

OpenStudy (lgbasallote):

\[\huge x^x + x^x = 839\] \[\huge \implies 2x^x = 839\] \[\huge \implies x^x = \frac{839}{2}\] \[\huge \implies x \ln x = \ln (\frac{839}{2})\] i got this far...

OpenStudy (anonymous):

 maybe next take log of both sides